cassandra has a total of 53 dimes and quarters. the dimes have a value of $6.60 less than the quarters. how much money does she have
Answer by ankor@dixie-net.com(22740)

Let d = no. of dimes
Let q = no. of quarters
:
Write an equation for each statement:
:
"Cassandra has a total of 53 dimes and quarters."
d + q = 53
d = (53-q); use this for substitution
:
"the dimes have a value of $6.60 less than the quarters."
.25q - .10d = 6.60
replace d with (53-q)
.25q - .10(53-q) = 6.60
.25q - 5.30 + .10q = 6.60
.25q + .10q = 6.60 + 5.30
.35q = 11.90
q = 11.9/.35
q = 34 quarters
then
53 - 34 = 10 dimes
Find the total value
.25(34) + .10(19) = $10.40
:
:
You can confirm these solutions: .25(34) - .10(19) = 6.60